Suppose that Head-First Company now sells both bicycle helmets and motorcycle helmets. The bicycle helmets are priced at $76 and have variable costs of $43 each. The motorcycle helmets are priced at $205 and have variable costs of $130 each. Total fixed cost for Head-First as a whole equals $59,850 (includes all fixed factory overhead and fixed selling and administrative expense). Next year, Head-First expects to sell 5,250 bicycle helmets and 2,100 motorcycle helmets.

Required:
1. Form a package of bicycle and motorcycle helmets based on the sales mix expected for the coming year.
2. Calculate the break-even point in units for bicycle helmets and for motorcycle helmets.
3. Check your answer by preparing a contribution margin income statement.

Amount Descriptions

Refer to the list below for the exact wording of text items within your income statement.

Amount Descriptions
Operating income
Operating loss
Sales
Total contribution margin
Total fixed cost
Total variable cost

Sales Mix and Break-Even Point

1. Form a package of bicycle and motorcycle helmets based on the sales mix expected for the coming year.

Product Price Unit Variable Cost Unit Contribution Margin Sales Mix Package Contribution Margin
Bicycle helmet
Motorcycle helmet
Package total

2. Calculate the break-even point in units for bicycle helmets and for motorcycle helmets.

Break-Even Bicycle Helmets
Break-Even Motorcycle Helmets

Contribution Margin Income Statement

3. Check your answer by preparing a contribution margin income statement. Refer to the list of Amount Descriptions for the exact wording of text items within your income statement.

Head-First Company

Contribution Margin Income Statement

At Break-Even Point

1

2

3

4

5
Answer #1
Part-1: Formulation of Package of Bicycle and Motorcycle helmeet- Head First company
Product Price Unit Variable Cost Unit Contribution Margin Sales Mix Package Contribution Margin
Bicycle helmet 76 43 33 0.7143 $23.6
Motorcycle helmet 205 130 75 0.2857 $21.4
Package total $45.0
*Sales Mix= Bicycle = 5250/7350, Motorcycle= 2100/7350
Computation of BEP
Bicycle helmet (1330*0.7143) 950
Motorcycle helmet (1330*0.2857) 380
*(Fixed Cost/Contribution Margin) ( 59850/45)                  1,330.00
Statmment of income statement
Bicycle helmet Motorcycle helmet Total
  Sales (950*76), ( 380*205) $72,201 $77,896
  Variable expenses $40,851 $49,398
  Contribution margin $31,351 $28,499 $59,849
  Fixed expenses $59,850
  Net operating profit ( will be 0 but -1 due to rounded off) -$1